OpenStudy (anonymous):

Simplify the given equation. 17x - 6 + 3x - 5 = x + 11 + 4x 20x - 11 = 5x + 11 20x + 1 = 5x + 11 9x = 16x

OpenStudy (anonymous):

20x - 11 = 5x + 11 <- you were right until here. 20x -11 + 11 = 5x + 11 + 11 20x = 5x + 22 20x - 5x = 5x - 5x + 22 15x = 22 x = 22/15

Parth (parthkohli):

-> 20x - 11 = 5x + 11

Parth (parthkohli):

First one is the correct answer